OpenStudy (anonymous):

Sara bought 8 fish. Every month the number of fish she has doubles. After M months she will have F fish, where F=8(2)^m How many fish will Sara have after 2 months if she keeps all of them and the fish stay healthy?

OpenStudy (freckles):

this is a plug in type of question

OpenStudy (freckles):

m represents the amount of months and F represents the amount of fish after m months

OpenStudy (freckles):

you want to know F when m=2

OpenStudy (anonymous):

Polygon ABCD is reflected and dilated to give polygon PQRS. The coordinates of the preimage are (2, 2), (6, 8), (12, 8), and (16, 2). The coordinates of the image are (11, 15), (9, 12), (6, 12), and (4, 15). What is the scale factor of the dilation?

OpenStudy (anonymous):

F=8(2)^2 F= 8*2^2 F= 8*4 F= ?

OpenStudy (anonymous):

what is this anser

OpenStudy (anonymous):

What is 8*4? That should be your answer.

OpenStudy (anonymous):

no this one

OpenStudy (anonymous):

Polygon ABCD is reflected and dilated to give polygon PQRS. The coordinates of the preimage are (2, 2), (6, 8), (12, 8), and (16, 2). The coordinates of the image are (11, 15), (9, 12), (6, 12), and (4, 15). What is the scale factor of the dilation?

OpenStudy (anonymous):

@brandon1919 please dont comment your questions on my question

OpenStudy (anonymous):

Nester; Did you get the answer you were looking for

OpenStudy (anonymous):

Yea i believe i did thank you

OpenStudy (anonymous):

i need help what mines

OpenStudy (anonymous):

@brandon1919 Post a separate question to get help on yours.

OpenStudy (anonymous):

i did do u know it